OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve f[g(6)] and select the correct answer below: f(x) = 6x + 12 g(x) = x − 8

OpenStudy (mathstudent55):

First, find g(6). Then use that value in function f.

OpenStudy (mathstudent55):

What is function g evaluated at x = 6? g(x) = x - 8 g(6) = 6 - 8 g(6) = ?

OpenStudy (anonymous):

-2 ?

OpenStudy (anonymous):

So , the answer would be 0 because 6(-2)+12=0 Right ?

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

Thank you so much ! Can I ask another one ?

OpenStudy (mathstudent55):

f(g(x)) = f(x - 8) = 6(x - 8) + 12 f(g(2)) = 6(2 - 8) + 12 = 6(-6) + 12 = -12 + 12 = 0

OpenStudy (mathstudent55):

You're welcome. Yes, go ahead.

OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve f over g(−4) and select the correct answer below: f(x) = 4x − 4 g(x) = x − 1

OpenStudy (mathstudent55):

\(\dfrac{f}{g}(x) = \dfrac{f(x)}{g(x)} = \dfrac{4x - 4}{x - 1} = \dfrac{4(x - 1)}{x - 1} = 4\) Since \(\dfrac{f}{g}(x) = 4\), no matter what x is, \(\dfrac{f}{g}(-4) = 4\)

OpenStudy (anonymous):

So technically x=4 ?

OpenStudy (mathstudent55):

No. In this problem, you were asked to find the value of the quotient of the functions at x = -4. For any value of x, the value of the quotient of the functions is 4, with only one exception, which is at x = 1. This division of functions is undefined at x = 1 because it involves division by zero. For all other values of x, the division of these two functions is 4.

OpenStudy (usukidoll):

@mathstudent55 factored a 4 out and was able to cancel x-1 so we don't have to evaluate when x = -4 anymore we just have 4

OpenStudy (usukidoll):